OpenStudy (anonymous):

Medal anyone....? A monk has a very specific ritual for climbing up the steps to the temple. First he climbs up to the middle step and meditates for 1 minute. Then he climbs up 8 steps and faces east until he hears a bird singing. Then he walks down 12 steps and picks up a pebble. He takes one step up and tosses the pebble over his left shoulder. Now, he walks up the remaining steps three at a time which only takes him 9 paces. How many steps are there? :))

OpenStudy (anonymous):

x=2(3*9+1-12+8)+1 = 2(24)+1 = 49 steps.